OpenStudy (anonymous):

The Shinkansen, Japan's high-speed trains, have been in service sine 1964. Sine that time, several train designs have been developed. Most of these trains travel between 240 km/h and 285 km/h. The exceptions are the "0" series, which began service in 1964, and the "500" series, which began service in 1997. Series 0 trains travel up to 220. km/h and have a total mass of about 8.84 x 10 kg. The lighter, streamlined series 500 trains travel up to 320.0 km/h, and have an estimated total mass of about 4.80 x 10 kg. What are the maximum kinetic energies that can be achieved by each of these trains

OpenStudy (anonymous):

\[ \text{kinetic energy} \ = \ \frac{1}{2}mv^2\] To have units of joules, remember to covert the velocities into m/s \[\frac{1 \ km}{h} = \frac{1000 \ m}{3600 \ s} = \frac{5 \ m}{18 \ s}\]

OpenStudy (anonymous):

"The Shinkansen, Japan's high-speed trains, have been in service sine 1964. Sine that time, several train designs have been developed. Most of these trains travel between 240 km/h and 285 km/h. The exceptions are the "0" series, which began service in 1964, and the "500" series, which began service in 1997." That's all useless information "Series 0 trains travel up to 220. km/h and have a total mass of about 8.84 x 10 kg. The lighter, streamlined series 500 trains travel up to 320.0 km/h, and have an estimated total mass of about 4.80 x 10 kg. What are the maximum kinetic energies that can be achieved by each of these trains" This is what you need to solve the problem.